OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 8x + 5y = 3 7x – 3y = 10

OpenStudy (shamil98):

First multiply both equations by a number to either cancel out x or y.

OpenStudy (shamil98):

In this case you would want to cancel out y, and then solve for x .

OpenStudy (ranga):

Part 1: To eliminate x: Multiply each term of the first equation by the coefficient of x in the second equation. Multiply each term of the second equation by the coefficient of x in the first equation. Subtract them to get rid of x and solve for y. Then solve for x. For part 2 use the method outlined in Part 1 to solve for x and y.

OpenStudy (anonymous):

i need like specific stuff.. ive tried and got it wrong every time...

OpenStudy (ranga):

8x + 5y = 3 7x - 3y = 10 Multiply each term of first equation by 7: 56x + 35y = 21 Multiply each term of second equation by 8: 56x - 24y = 80 Subtract the above two equations: (56x - 56x) + (35y - (-24y)) = (21 - 80) 59y = -59 So y = -1 Put y = -1 in 8x + 5y = 3 and solve for x

OpenStudy (anonymous):

thank you! that helped out alot !